OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#8945 — Optical loop Seine-Saint-Denis (93)
Incident Report for Web Cloud
Resolved
There is a cut on the NRA optical loop in departments 93 and 94, securisation is playing its part and there has not been any downtime.

We have opened a ticket with the fibre optic operator so that they intervene on the stretch of fibre affected.

Update(s):

Date: 2013-07-12 05:31:34 UTC
The fibers were remounted. Everything is functional.

Date: 2013-07-12 00:42:16 UTC
Our fibers have not been properly re-welded due to a reversal in the fiber tube labeling .
The welding is in progress.
It takes 3 hours.

Date: 2013-07-12 00:40:16 UTC
A problem on the other side of the loop isolated STO93 completely.
A team goes on site.

Date: 2013-07-12 00:39:28 UTC
The cut-off is due to civil works, the optical fiber was pulled over a few meters.

Date: 2013-07-11 16:09:41 UTC
The intervention is underway, the technicians are currently on the optical junction box searching for the affected point.

Date: 2013-07-11 13:49:49 UTC
The cut was located 2100 metres from the STO93 site. The team is moving to this location.

Date: 2013-07-11 11:09:40 UTC
An intervention has been launched to test the fibre optic on site. A team has been deployed.
Posted Jul 11, 2013 - 09:14 UTC